OpenStudy (anonymous):

Find the linear approximation of f(x)=ln x at x=1 and use it to estimate ln 1.16. L(x)= . ln 1.16 approx

OpenStudy (freckles):

do you know how to find the tangent line to the curve f(x)=ln(x) at x=1?

OpenStudy (anonymous):

no, i'm sorry

OpenStudy (freckles):

The tangent line to the curve y=f(x) at x=a is \[y-f(a)=f'(a) \cdot (x-a)\]

OpenStudy (freckles):

does that look familiar?

OpenStudy (freckles):

If f(x)=ln(x) then f'(x)?

OpenStudy (freckles):

f'(x)=?

OpenStudy (anonymous):

1/x?

OpenStudy (freckles):

good now what is f'(1)

OpenStudy (freckles):

if f'(x)=1/x then f'(1)=1/1 right? so the tangent line is \[y-f(1)=f'(1)(x-1) \\ y-\ln(1)=\frac{1}{1}(x-1)\]

OpenStudy (freckles):

I will let you finish simplifying that

OpenStudy (anonymous):

\[\frac{1}{1}\]

OpenStudy (freckles):

\[y-f(a)=f'(a) \cdot (x-a) \\ y=f'(a)(x-a)+f(a) \\ f(x) \approx f'(a)(x-a)+f(a) \text{ for values near } x=a\]

OpenStudy (anonymous):

you plug it in for a?

OpenStudy (anonymous):

i'm sorry if I misunderstood

OpenStudy (anonymous):

no \(a=1\) replace \(x\) by \(1.16\)

OpenStudy (anonymous):

I still understand i'm so sorry

OpenStudy (anonymous):

Dont*

OpenStudy (freckles):

what part is hard to understand

OpenStudy (anonymous):

of finding ln1.16 approx

OpenStudy (freckles):

\[f(1.16) \approx f'(1)(1.16-1)+f(1)\]

OpenStudy (freckles):

we already found f'(1) and f(1)

OpenStudy (anonymous):

fprime(1) is 1 and f(1) is ln(1) right?

OpenStudy (freckles):

yes but ln(1)=?

OpenStudy (anonymous):

0 i believe

OpenStudy (freckles):

yes

OpenStudy (anonymous):

so its 0.16?

OpenStudy (freckles):

so you have f(1.16) is approx 1.16-1

OpenStudy (freckles):

yes

OpenStudy (anonymous):

then l(x)= the same thing?

OpenStudy (freckles):

what is l(x)?

OpenStudy (freckles):

oh is that 1*x 1*x is x

OpenStudy (freckles):

i don't know what you meant about it equaling the same thing though

OpenStudy (anonymous):

oh i thought the answer was the same because it 2 separate questions

OpenStudy (freckles):

i don't understand what we are talking about anymore

OpenStudy (anonymous):

lmao i'm sorry, if you look back to the question, they're asking for 2 separate answers

OpenStudy (anonymous):

their asking what is L(x) and what is ln1.16

OpenStudy (freckles):

We already found the tangent line to x=a and then used it to approximate ln(1.16)

OpenStudy (freckles):

tangent line to f at x=a*

OpenStudy (freckles):

our a was 1

OpenStudy (anonymous):

I'm sorry, so what is l(x) you probably stated this in your question

OpenStudy (freckles):

\[y-f(a)=f'(a)(x-a) \\ y=f'(a)(x-a)+f(a) \\ L(x)=f'(a)(x-a)+f(a) \\ f(x) \approx L(x) \\ f(x) \approx f'(a)(x-a)+f(a)\]

OpenStudy (freckles):

L(x) is f'(a)(x-a)+f(a)

OpenStudy (freckles):

but this was already found above

OpenStudy (freckles):

remember f'(x)=1/x f'(1)=1 f(1)=ln(1)=0

OpenStudy (anonymous):

and what is a

OpenStudy (anonymous):

nvm i know a

OpenStudy (freckles):

\[L(x)=(\frac{1}{x})_{x=1}(x-1)+\ln(1)\]

OpenStudy (freckles):

(1/x) evaluated at x=1 is of course 1/1=?

OpenStudy (anonymous):

its x-1

OpenStudy (freckles):

yes

OpenStudy (anonymous):

Omg

OpenStudy (freckles):

lol

OpenStudy (anonymous):

I'm so sorry i feel like i frustrated you so much

OpenStudy (freckles):

like that is what we plugged 1.16 into

OpenStudy (anonymous):

Thank you for all your help. Again i'm sorry

OpenStudy (freckles):

to approximate ln(1.16)

OpenStudy (freckles):

np

OpenStudy (anonymous):

I appreciate it a lot, i know it's frustrating to teach me

OpenStudy (freckles):

nah not really

OpenStudy (anonymous):

Thanks :)